# Current distribution and chirality in the regime of the quantized Hall effect

### v1

@@ Abstract
The integer quantum Hall effect (QHE) is one of the most basic and, at the same time, astonishing phenomena in condensed matter physics. Despite its importance for metrology and fundamental physics, our microscopic understanding of the QHE is not yet complete. For decades, there has been a discussion about the current distribution inside a Hall bar for the case of a quantized Hall resistance. With the advent of quantum technology, the questions, if the current flows in compressible Landauer-Büttiker edge channels or in incompressible strips with varying geometry and whether the current flow is chiral or not, gain importance, because the quantum phase of each charge carrier depends on its exact pathway. In the present work, we compare multi-terminal current measurements in the regime of the QHE with model calculations in order to unravel the current density distribution in a Hall-bar device. We find, that the conjecture of chirality is justified for the plateaus of quantized Hall resistance, while in between plateaus the current density looses its chirality. The latter is described by the Drude model, which predicts a homogeneous current density for a sample with homogeneous mobility and carrier density.
@@ I Introduction
Magnetic field is an axial vector and, therefore, breaks time-reversal symmetry, which, however, can be restored by the simultaneous reversal of magnetic field and time, $\vec{B}\to-\vec{B}$ and $t\to-t$. This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$. The axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ in the classical equation of motion $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$ of an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$. As an introduction, we discuss the dynamics of charge carriers considering a variety of boundary conditions.
@@
The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a constant drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf. Fig. 1(a). The chirality of this trajectory is left-handed for the negatively charged electron (but would be right handed for a positive charge). Note that we define the chirality of a trajectory in analogy to the handedness of a screw thread.
@@
Figure 1: Motion (expressed by drift velocity) of an electron confined to the $xy$-plane with in-plane electric and perpendicular magnetic fields, $\vec{E}$ and $\vec{B}$ (red arrows), respectively. (a) Ballistic limit (free electron). The spiral is the classical trajectory consisting of a constant drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity. (b) Diffusive limit. For $|\mu B|<1$ the spiral motion is damped. However, the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ and tilts it by the Hall angle $\varphi$. (c) Diffusive Hall bar in the steady state limit with edges parallel to the applied $\vec{E}$ (red arrow). The Hall voltage compensates the Lorentz force, $-e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$ (or $E_{y}=|\vec{F}_{\text{L}}|/e$), such that the drift velocity is $\vec{\upsilon}_{\text{d}}$ as for $\vec{B}=0$.
@@
The Drude model describes magnetotransport in the classical limit of the diffusive regime. Its key parameter is the electron mobility $\mu$. For $\mu B<1$ (corresponding to a mean-free path $l_{\text{m}}<R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering. Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$ and applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$, where $\tau=m\mu/e$ is the momentum scattering time. In this diffusive limit and for $\vec{B}\perp\vec{E}$, the solution of the equation of motion introduced above can be written as
$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,.$ (1)
@@
[For $\vec{E}\not\perp\vec{B}$ there would be a third summand, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$, inside the parantheses.] Clearly, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous. However for finite $B$, $\vec{j}$ is not parallel to $\vec{E}$, but bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$, cf. Fig. 1(b).
The ballistic limit discussed above is clearly beyond the scope of the Drude model. Indeed, for $\mu B\to\infty$, the friction forces in Eq. (1) become negligible compared to the Lorentz force and the drift velocity becomes independent of $\mu$. Nevertheless for $\mu B\to\infty$, the drift velocity takes the value $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, which is identical to that of a ballistic electron. This observation encourages us to compare measurements with the Drude model even for $\mu B\gg 1$.
@@
In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, and include boundaries of the 2DES and electron scattering. Furthermore, we consider an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular, homogeneous magnetic field $\vec{B}=(0,0,B)$, corresponding to one of the most common set-ups for magnetotransport measurements. For simplicity, we restrict our general discussion to a homogeneous electron density and mobilty inside the 2DES (away from the edges).
@@
To discuss the classical limit of the Hall effect, we include two sample edges at $y=0$ and $y=b$, such that the electrons are confined in the $y$-direction but move freely in the $x$-direction, cf. Fig. 1(c). In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons. Applying this condition to Eq. (1) we find $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$. The transversal field $E_{y}=\vec{\nabla}V_{\text{H}}$ is maintained by charge accumulation at the sample edges, which gives rise to the Hall voltage $V_{\text{H}}$. In the steady state, $-eE_{y}=-F_{\text{L}}$, i.e., the Lorentz force acting on a free electron is exactly canceled, and the electrons move as if no magnetic field was applied but only the longitudinal field $E_{x}$. For our homogeneous sample, the current density $j_{x}$ is a constant and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
In a confined sample, the axial character of the magnetic field manifests itself in the edge charges or the related Hall voltage (which break the time reversal symmetry). The latter gives rise to an increase of the two-terminal resistance of the 2DES according to $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}$ with the resistance $R_{0}$ of the 2DES at $B=0$.
@@
As we increase $B$, the quantization of the density of states into Landau-levels results in the quantized Hall effect (QHE). Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance. The latter indicates scattering-free transport within the Hall bar.
@@
To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of perfect one-dimensional (1D) edge channels, in which scattering is suppressed. The conductance of a spin-resolved 1D channel is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$. The filling factor $\nu$ describes the number of perfect 1D channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk. The perfect 1D-channels are assumed to follow the edges of the Hall bar, because they form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2]. To obtain a quantized Hall resistance, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization. The direction of the current flow in the edge channels is predetermined by the gradient of the confinement potential. This yields current flow in opposite directions at opposite edges. In that sense, in the LBP the QHE is associated with chiral current flow.
@@
The LBP provides an intuitive explanation for the quantized Hall resistance and the vanishing longitudinal resistance of the QHE. While its merit is its simplicity, its predictions are limited due to the phenomenological character as well as its neglect of interactions and non-equilibrium effects. In particular, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [3, 4]. However, this would cause highly charged stripes between the edge channels corresponding to an unrealistically large Coulomb energy.
The screening theory [3, 4, 5, 6, 7, 8, 9] overcomes these limitations by taking into account the direct Coulomb interaction between the electrons. The Coulomb interaction causes a rearrangement of the free carriers such that the highly charged stripes are avoided. Instead, it leads to a segmentation of the Hall bar in compressible regions and incompressible strips (ICSs). Inside the ICSs, carriers cannot scatter and screening of the electric field is absent, because the chemical potential is in a local energy gap between filled and empty Landau levels; ICSs are accompanied by adjacent compressible regions, in which a Landau-level is aligned with the chemical potential, such that carriers can scatter. Inside the compressible regions, the electric field is then perfectly screened. The current flows where the electric field drops, i.e., inside the ICSs. In contrast to the 1D Landauer-Büttiker edge channels, the ICSs have a finite width exceeding the magnetic length [7, 8, 9]. Width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$. In particular, towards the end of each plateau with increasing magnetic field, the screening theory predicts a single extended incompressible region in the center of the Hall bar [7, 8, 9]. This implies that current can flow in the bulk of the Hall bar and not just in edge channels. The predicted bulk ICS near the high magnetic field end of the plateaus was experimentally corroborated in scanning gate spectroscopy experiments [10, 11, 12] and recently directly confirmed in transport experiments [13]. As we will discuss in Sec. IV below, the screening theory is consistent with chiral transport at the plateaus of the QHE, even if the current flows inside a single bulk ICS.
In the present article, we investigate the influence of a magnetic field on the current density distribution of a multiterminal junction. In Sec. II we illuminate the non-chiral character of the Drude model and derive a general coherent and chiral model. In Sec. III we compare our multi-terminal current measurements with the predictions of the models. For relatively small magnetic fields, we find that the classical limit of the Hall effect is described by the Drude model. For higher magnetic fields, we distinguish between the plateau regions of a quantized Hall resistance and the transition regions between plateaus. For the plateaus, coherent and chiral transport was already supported by Mach-Zehnder interferometer experiments measuring interference between two distinct current paths [14]. Our own results confirm this. For the magnetic field regions in between the plateaus our measurements are consistent with the predictions of the classical Drude model, which predicts non-chiral transport. In Sec. IV we explain, why the screening theory predicts chiral transport.
II model
Figure 2: Drude model: Distribution of current $I=I_{xx}+I_{xy}$ in a three-terminal device with homogeneous electron density and mobility in the diffusive and classical limit for $V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{2}=R_{3}$ and $\vec{B}$ pointing into the plane of the 2DES. Color intensities indicate the distribution of the absolute value of the current density into the three branches. Within each branch, the current density is homogeneous. For $V_{1}<0$, white arrows show the directions of the conventional currents and black bent arrows the direction and relative magnitudes of electron flow (for $V_{1}>0$ vice versa). (a) $I_{xx}$ displaying ohmic behavior, the same as for $B=0$: $I_{xx}^{(2)}=I_{xx}^{(3)}=-0.5I_{xx}^{(1)}$. (b) $I_{xy}$ for $R_{2}=R_{3}=0$ displaying non-ohmic behavior: $I_{xy}$ is bent by the Lorentz force: $I_{xy}^{(2)}=-I_{xy}^{(1)}$ and $I_{xy}^{(3)}=0$. (c) $I_{xy}$ for $R_{2}=R_{3}>0$: $I_{xy}$ is still bent by the Lorentz force but partly reflected at the resistive contacts: $I_{xy}^{(2)}>I_{xy}^{(3)}$ with $I_{xx}^{(2)}+I_{xx}^{(3)}=-I_{xx}^{(1)}$. The distribution of $I_{xy}$ resembles chiral behavior but $I_{xy}$ is homogeneous in each branch (away from the junction region).
In this section, we discuss the Drude model and, using the Landauer-Büttiker formalism (not to be confused with the LBP), derive a coherent and chiral model.
The magnetic field regions in between quantized plateaus deserve special attention. Because, near the plateaus of the QHE, the quantization of the density of states into Landau levels is relevant, we can classify these regions as belonging to the quantum regime. However, between plateaus, the electron transport is diffusive and our coherent and chiral model fails. In fact, the mobility $\mu(B)$ strongly oscillates, such that $\mu B$ can be in the order of 1, as determined from the Shubnikov-de-Haas oscillations of the longitudinal resistance. Can we then apply the Drude model at large magnetic fields for the regions between quantized plateaus and is the magnetotransport between the plateaus chiral or not? In Sec. III we will provide an answer to these questions.
II.1 Drude model for a three-terminal junction
Using $\vec{j}=en_{\text{s}}\vec{\upsilon}_{\text{d}}$ we can rewrite Eq. (1) in terms of the magnetic-field-dependent conductivity tensor as
where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$. For a laterally confined 2DES in the steady state, we assume $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$. Following the structure of the conductivity tensor, we devide the current density $j_{x}$ into two components
$j_{x}\equiv j_{xx}+j_{xy}\quad\text{with}$
$j_{xx}=\frac{en_{\text{s}}\mu}{1+(\mu B)^{2}}\,E_{x}\quad\text{and}\quad j_{xy}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{\text{H}}}\,E_{y}\,.$
The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is driven by the field $E_{x}$, which acts along the Hall bar because of its finite resistivity $\sigma_{0}^{-1}$. The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is linked to the Hall voltage and attenuated by the Hall resistance. For $B=0$ the component $j_{xy}=0$, while for $\mu B\to\infty$ the current density becomes independent of $\mu$ as $j_{xx}\to 0$.
[…]
$I=I_{xx}+I_{xy}\quad\text{with}$ (4)
$\frac{I_{xx}}{I}=\frac{1}{1+(\mu B)^{2}}\quad\text{and}\quad\frac{I_{xy}}{I}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,.$
[…]
First, we discuss the current components $I_{xx}$ and $I_{xy}$ for the relatively simple case of a symmetric three-terminal junction as sketched in Fig. 2. We apply a voltage $V_{1}$ to contact 1 while the other two contacts are connected to the electrical ground, $V_{2}=V_{3}=0$. Each contact has an ohmic contact resistance $R_{1,2,3}$ and we apply a finite magnetic field perpendicular to the conducting plane and pointing into it. We can measure the currents $I_{1,2,3}=I_{xx}^{(1,2,3)}+I_{xy}^{(1,2,3)}$. They fulfill Kirchhoffs junction rule $I_{1}+I_{2}+I_{3}=0$ also for the components $I_{xx}^{(1)}+I_{xx}^{(2)}+I_{xx}^{(3)}=0$, $I_{xy}^{(1)}+I_{xy}^{(2)}+I_{xy}^{(3)}=0$, because the respective driving fields are prependicular to each other. The longitudinal component $-I_{xx}^{(1)}$ will be divided between contacts 2 and 3 according to Ohm’s law as indicated in Fig. 2(a): $I_{xx}^{(2,3)}=-I_{xx}^{(1)}\frac{R_{3,2}}{R_{2}+R_{3}}$. The transversal component $-I_{xy}^{(1)}$ follows the charged edge of the Hall bar and is accordingly bent into contact 2 (or into contact 3 for $\vec{B}$ pointing out of the plane). In Fig. 2(b), we indicate the transversal component for the simplified case of a vanishing contact resistance, $R_{2}=0$, resulting in $I_{xy}^{(2)}=-I_{xy}^{(1)}$ and $I_{xy}^{(3)}=0$.
In a realistic device with finite contact resistances, the voltage drop across each contact resistance causes a partial reflection at each contact. Therefore, a smaller fraction of $-I_{xy}^{(1)}$ flows from contact 2 to contact 3 as shown in Fig. 2(c). Because the magnitude of $I_{xy}^{(1)}$ is predetermined by the applied voltage and the Hall resistance, we can express the division of $-I_{xy}^{(1)}$ in terms of two parallel channels with the effective resistances $R_{2}$ versus $R_{\text{H}}+R_{3}$. Hence, we have $I_{xy}^{(2)}=-I_{xy}^{(1)}\frac{R_{\text{H}}+R_{3}}{R_{\text{H}}+R_{2}+R_{3}}$ and $I_{xy}^{(3)}=-I_{xy}^{(1)}\frac{R_{2}}{R_{\text{H}}+R_{2}+R_{3}}$.
The ohmic components $I_{xx}^{(1,2,3)}$ vanish quickly for $\mu B\gg 1$. The transveral currents $I_{xy}^{(1,2,3)}$ vanish at $B=0$, but for $\mu B\gg 1$ the current density distribution merits a closer look. On the one hand, away from a junction, the Drude model predicts a homogeneous current density across a homogeneous sample, cf. Eq. (1), classifying $j_{x}$ to be non-chiral. On the other hand, near a junction, the edge charges caused by the Lorentz force break the homogeneity of the transversal current density $j_{xy}$, while the current $I_{xy}$ is bent into just one of the branches of the junction. Finally, if considering the currents flowing out of each contact in a measurement at $\mu B\gg 1$, practically being a measurement of the transversal components $I_{xy}^{(1,2,3)}$, we expect to find identical result whether the transport is chiral or not. Hence, an experimental indication of non-chiral transport in a multiterminal current measurement can only be expected for $\mu B\lesssim 1$, where the homogeneous $j_{xx}$ is sizable.
II.2 Coherent and chiral model of $n$-terminal device

### v2

# Chiral nature of current flow in the regime of the quantized Hall effect

@@ Abstract
The integer quantum Hall effect (QHE) belongs to the most fundamental phenomena of solid state physics and has an important application as resistance standard. It serves as a basis to understand the fractional, anomalous or spin QHEs, candidates for applications in quantum technology due to their topological properties. For optimizing all these applications it is essential to understand the current distribution inside the Hall bar, a question disputed for decades. We perform multiterminal current measurements on a Hall bar and compare the results with limiting models. We show, based on these simple experiments, that the current flow is chiral for the plateaus of quantized Hall resistance. Everywhere else, the current is described by the Drude model.
@@ I Introduction
Magnetic field is an axial vector and, therefore, breaks time-reversal symmetry, which, however, can be restored by the simultaneous reversal of magnetic field and time, $\vec{B}\to-\vec{B}$ and $t\to-t$. This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$. In this article, we explore the implication of the axial symmetry of $\vec{B}$ for the current flow in the regime of the quantized Hall effect.
In the classical equation of motion, $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$, the axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ acting on an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$.
@@
The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf. Fig. 1(a). If we define the chirality of a trajectory in analogy to the handedness of a screw thread, the above trajectory of a ballistic electron is left-handed (and would be right handed for a positive charge).
@@
Figure 1: Motion (expressed by drift velocity) of an electron confined to the $xy$-plane with in-plane electric and perpendicular magnetic fields, $\vec{E}$ and $\vec{B}$ (red arrows), respectively. (a) Ballistic limit (free electron). The spiral is the classical trajectory consisting of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity. (b) Diffusive limit. For $|\mu B|<1$ the spiral motion is damped. However, the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ and tilts it by the Hall angle $\varphi$. (c) Diffusive Hall bar in the steady state limit with edges parallel to the applied $\vec{E}$ (red arrow). The Hall voltage compensates the Lorentz force, $-e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$ (or $E_{y}=|\vec{F}_{\text{L}}|/e$), such that the drift velocity is $\vec{\upsilon}_{\text{d}}$ as for $\vec{B}=0$.
@@
The Drude model describes magnetotransport in the classical limit of the diffusive regime. Its key parameter is the electron mobility $\mu$. For $\mu B\lesssim 1$ (corresponding to a mean-free path $l_{\text{m}}\lesssim R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering. Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$. Applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$ with the momentum scattering time $\tau=m\mu/e$, the solution of the equation of motion introduced above in the diffusive limit becomes
$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,\text{ for }\vec{B}\perp\vec{E}\,.$ (1)
@@
[For $\vec{B}\not\perp\vec{E}$ there would be a third summand inside the parantheses, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$.] For finite $B$, $\vec{j}$ is bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$ in respect to $\vec{E}$, cf. Fig. 1(b). Importantly, as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous, too.
While the Drude model is designed to describe the limit of $\mu B\lesssim 1$, its prediction for the drift velocity for $\mu B\to\infty$ where the friction forces in Eq. (1) become negligible, namely $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, is identical to the ballistic solution. Since for $\mu B\to\infty$ the rapid cyclotron motion averges out over reasonable length scales, both models predict identical current densities, too, for $\mu B\to\infty$.
@@
In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, with an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular homogeneous magnetic field $\vec{B}=(0,0,B)$. This is a common scenario realized in many mesoscopic magnetotransport measurements.
@@
To discuss the classical limit of the Hall effect, we include two sample edges at $y=0$ and $y=b$, such that the electrons are confined in the $y$-direction but move freely in the $x$-direction, cf. Fig. 1(c). In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons. Applied to Eq. (1) $j_{y}=0$ yields $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$. Physically, the transversal field $E_{y}$ is sourced by charge accumulated at the sample edges. With $E_{y}=\vec{\nabla}V_{\text{H}}$ it also gives rise to the Hall voltage $V_{\text{H}}$, that can be measured across the Hall bar. Because in steady state, the Lorentz force, acting on the carriers, is exactly canceled by $-eE_{y}$ (the forces associated with $B$ and $E_{y}$ cancel each other), inside the Hall bar the electrons move as if only the longitudinal electric field $E_{x}$ was applied. Given a homogeneous sample (homogeneous distributions of $n_{\text{s}}$ and $\mu$ away from the edges of the Hall bar), the current density $j_{x}$ is uniform and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
While the Drude model predicts a homogeneous current density for the classical Hall effect, the axial character of the magnetic field and the breaking of the time-reversal symmetry are manifest in the edge charges and the related Hall voltage. In addition, the Hall resistance is non-local as it does not cause any dissipation inside the Hall bar: At $B=0$ the two terminal current measurement of a Hall bar is the sum of the ohmic sample resistance and the resistance of the leads (including the ohmic contacts), $R_{\text{2T}}(0)=R_{0}+R_{\textrm{leads}}$. At finite $B$, the non-local Hall resistance enhances the resistance of the Hall bar, such that $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}+R_{\textrm{leads}}$.
@@
As we increase $B$, at low enough temperatures the quantization of the density of states into discrete Landau-levels results in the quantum Hall effect (QHE). Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance ($R_{0}=0$). The latter indicates the absence of momentum scattering within the Hall bar for a quantized Hall resistance, in agreement with the classical Drude prediction for $\mu B\to\infty$.
@@
To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of chiral, one-dimensional (1D) edge channels, such that backward scattering is suppressed [1, 2]. The conductance of a spin-resolved 1D channel without momentum scattering is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$. The filling factor $\nu$ describes the number of edge channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk. The LBP assumes that the edge channels form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2]. To obtain finite ranges of quantized Hall resistances, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization. The direction of the current flow in the edge channels is predetermined by the gradient of the confinement potential. This yields chiral current flow, more precisely, current flow in opposite directions along opposite edges.
@@
The LBP provides an intuitive explanation for the quantized Hall resistance and the vanishing longitudinal resistance of the QHE. While its merit is its simplicity, its predictions are limited due to the phenomenological character as well as its neglect of interactions and non-equilibrium effects. In particular, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [3, 4]. Hence, neglecting the Coulomb interaction between carriers in the LBP causes the formation of highly charged stripes between the edge channels. This local polarization corresponds to an unrealistically large Coulomb energy. Another weak point of the LBP is the coexistence of occupied and unoccupied states inside the current carrying edge channels at the chemical potential. Even, if we assume chiral edge channels, an almost complete suppression of carrier scattering, as indicated by the accuracy of the measurements of the von Klitzing constant, is unlikely under any realistic conditions.
The screening theory [3, 4, 5, 6, 7, 8, 9] offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$. By taking into account the direct Coulomb interaction between electrons, it overcomes the limitations of the LBP. The Coulomb interaction causes a rearrangement of the free carriers such that the highly charged stripes are avoided. Early purely classical formulations considered this screening based on the Thomas-Fermi approximation [3, 4]. They demonstrated that the Landau-quantization leads to a segmentation of the 2DES of a Hall bar in compressible regions and incompressible strips (ICSs). The term ICS denotes areas of the 2DES, where the chemical potential lies within a local energy gap between completely filled and empty Landau levels. For this reason, at low enough temperature, inside the ICSs carriers cannot scatter and screening of the electric field is absent. ICSs are accompanied by adjacent compressible regions, in which a partly filled Landau-level is aligned with the chemical potential. Hence, inside compressible regions, carriers can scatter. Provided the vicinity of unscreened ICSs, the screening of the local electric field becomes perfect inside the compressible regions. The early calculations already show, that both, width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$. Later, for given boundary conditions the number and geometry of the ICSs were predicted more accurately using a self-consistent numerical calculation of Poisson’s equation within the Thomas-Fermi approximation and accounting for the quantum mechanical wavefunctions of the electrons in a mean field approach [7, 8, 9] (later confirmed in a full self-consistent approach of the quantum- electrostatic problem [10]). The calculations predict, that towards the higher magnetic field end of each plateau the ICSs merge into a single extended incompressible region centered in the Hall bar [7, 8, 9]. These predictions clearly confirmed the electrostatic landscape measured in scanning probe spectroscopy experiments as a function of the magnetic field [11, 12, 13].
According to the screening theory the electric field vanishes inside the perfectly screened compressible regions while it is exclusively contained inside the ICSs. For suppressed scattering the drift velocity of the electrons is $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$. Hence, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, vanishes in the compressible regions, while all current flows inside the ICSs [4, 14, 9, 15]. In case of a bulk ICS, current should then flow in the center of the Hall bar. Recently, we confirmed the predicted bulk current directly by means of an additional tiny ohmic contact in the center of a Hall bar [16]. Both, the electrostatics and the existence of bulk currents, predicted by the screening theory [9] and confirmed in measurements [11, 12, 13, 16] are in direct contrast to the assumptions of the LBP. Therefore, it is important to find out, whether another assumption of the LBP, which is often used, remains correct, namely whether the current flow is chiral for the quantized plateaus of the Hall resistance.
Even within the screening theory this is still a controversal question [3, 4, 9, 10, 15]. In the present article, we explore the question of chirality experimentally by employing multiterminal current measurements. Thereby, we probe the division of the current into various ohmic contacts of the Hall bar. Clearly, we do not directly probe the local current density distribution. Nevertheless, it is possible to distinguish between chiral versus non-chiral transport as the current division into multiple contacts directly depends on it. To quantify this dependence, we use the properties of the macroscopic scattering matrix, which yields the division of the current into various contacts as follows: In the case of chiral current flow, only unidirectional transmission coefficients connecting adjacent contacts are non-zero. In comparison, non-chiral current flow requires additional non-zero transmission coefficients, which would ultimately yield a different divison of the current into the various contacts.
Here, we show that the measured current division can be described for arbitrary magnetic fields by either one of two limiting models, introduced in Sec. II below. In Sec. III, we then present our three- and four-terminal current experiments, compare them with the predictions of the two limiting models and discuss our results.
II models
Our first model generically assumes coherent and chiral current flow. The condition of coherence is technical. In practice, it is sufficient to assume the absence of scattering. The second model applies the classical Drude model to a multiterminal current measurement. It describes diffusive and non-chiral current.
Our experiments in Sec. III show, that the generic chiral model applies for the magnetic field regions of the plateaus of quantized Hall resistance, while the Drude model describes the situation for all other magnetic fields.
II.1 Coherent and chiral model of $n$-terminal device
Our first model describes the limit of coherent and chiral current flow inside the 2DES of a sample equipped with $n$ ohmic contacts, cf. Fig. 2.